OpenStudy (anonymous):

Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. (5 points) Part B: How many liters of orange juice are present in the can? Show your work. (5 points)

OpenStudy (kewlgeek555):

Well you know that 100% is one liter, and that there is 80% of orange juice, then you must assume there is 20%. So, I guess the equation could be 1 - 0.80 = 0.20 liters of water. And you know that there are 0.80 liters of orange juice. At least that's how I think it goes.

OpenStudy (anonymous):

i guess , but doesn't my equation need a variable? this word problem is confusing.

OpenStudy (anonymous):

Alright, to represent the variable as you require, it could be: 100% - 1l 80% - x Therefore x= 80*1/100, which stays as x=0.8 And, state volume of water as the variable 'y' Therefore y= 1litre - xliters Substitute the x with 0.8 and solve it out, getting 0.2 So your solution is: x=0.8 and y=0.2 Could this help?

OpenStudy (anonymous):

So for part a i would put x=80*1/100 ?

OpenStudy (anonymous):

I prenatal tha for you o get the 5 points, you should present in your paper the following lines which I wrote above ( consider the part of 'Alright, to represent the variable as you require. , it could be' as line 1 and 'could this help?' as line 10 ): So, you should put lines 2, 3, 4, 5, 6 and 9.

OpenStudy (anonymous):

thank you this helps :)

OpenStudy (anonymous):

You are most welcome.